OpenStudy (anonymous):

What equation can be used to represent the following sequence: 7,10,13,16

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

the pattern is , they go up by the same amount each time

OpenStudy (anonymous):

okay. thatd be 3,no?

OpenStudy (unklerhaukus):

right,

OpenStudy (anonymous):

7n+3?

OpenStudy (anonymous):

It is in nth term,no?

OpenStudy (anonymous):

o no nvm

OpenStudy (unklerhaukus):

we want the expression to equal seven when n=1

OpenStudy (anonymous):

okay. so, .

OpenStudy (anonymous):

lemme think

OpenStudy (unklerhaukus):

(because the first term have to be in the sequence)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

ok, brb. lemme try something

OpenStudy (unklerhaukus):

when n=1, 7n+3= 7+3 =10 so that's not right

OpenStudy (anonymous):

a(n)=3n+b a(1)=7 find b

OpenStudy (anonymous):

3n+4

OpenStudy (anonymous):

good

OpenStudy (anonymous):

thanks,

OpenStudy (anonymous):

np